# Near-Field Integrated Sensing and Communications

### v3

@@ Abstract
Index Terms:
Integrated sensing and communications (ISAC), joint distance and angle estimation, near-field.
@@ I Introduction
Furthermore, the near-field effect also provides new possibilities. The spherical wave propagation in the near field introduces a new distance dimension, which has the potential to facilitate joint estimation of distance [9] and angle in sensing and to mitigate interference in communications [10]. To the best of the authors’ knowledge, near-field ISAC systems have not been studied yet, which motivates this work. In this letter, we propose a near-field ISAC framework. Based on the spherical wave propagation, we establish the accurate near-field channel model of the near-field ISAC framework. We jointly optimize the ISAC signal to maximize the near-field sensing performance subject to the minimum near-field communication rate for both fully digital antennas and hybrid digital and analog antennas. Finally, our numerical results verify the effectiveness of the proposed framework.
@@ II System Model
As shown in Fig. 1, we consider a narrowband near-field ISAC system comprising an $N$-antenna dual-functional base station (BS), where $N=2\tilde{N}+1$, $K$ single antenna communication users, and one sensing target. We consider a monostatic sensing setup at the BS. Furthermore, we assume that the BS employs a uniform linear array (ULA) with an antenna spacing of $d$, resulting in an aperture of $D=(N-1)d$. Typically, the boundary between near-field and far-field can be determined by the Rayleigh distance $\frac{2D^{2}}{\lambda}$ [3], where $\lambda$ is the signal wavelength. We assume that the communication users and sensing target are located in the near-field region of the BS, which implies that their distance from the BS is less than $\frac{2D^{2}}{\lambda}$. As observed, in order to take advantage of the new features of the near field, ELAAs and high frequencies are required to produce a large near-field region. Therefore, in the following, we first discuss the optimal full digital antennas to provide a performance upper bound of the near-field ISAC system. Then, we extend to the hybrid digital and analog antennas, which are more energy-efficient for ELAAs with high frequencies.
@@ II-A Near-field Channel Model for ISAC
$[\mathbf{a}(r,\theta)]_{n}=e^{-j\frac{2\pi}{\lambda}(r_{n}(r,\theta)-r)}.$ (4)
@@
$[\mathbf{a}_{\mathrm{far}}(\theta)]_{n}=e^{-j\frac{2\pi}{\lambda}(-nd\cos\theta)}.$ (5)
@@ II-B ISAC Model for Fully Digital Antennas
$\mathbf{x}[t]=\sum_{k\in\mathcal{K}}\mathbf{f}_{k}c_{k}[t]+\mathbf{s}[t],\vspace{-0.2cm}$ (8)
where $\mathbf{f}_{k}\in\mathbb{C}^{N\times 1}$ denotes the fully digital beamformer for conveying the information symbol $c_{k}[t]\in\mathbb{C}$ to the user $k$, $\mathbf{s}[t]$ denotes the dedicated sensing signal for achieving the full sensing degrees of freedom [6], and $\mathcal{K}=\{1,\dots,K\}$. The information symbols are assumed to be independently distributed and have unit power, i.e., $\mathbb{E}\big{[}c_{k}[t]c_{i}^{*}[t]\big{]}=1$, if $k=i$; and $\mathbb{E}\big{[}c_{k}[t]c_{i}^{*}[t]\big{]}=0$, otherwise. Let $\mathbf{R}_{s}=\mathbb{E}\big{[}\mathbf{s}[t]\mathbf{s}^{H}[t]\big{]}$ denote the covariance matrix of the dedicated sensing signal. Then, the covariance matrix of the sensing signal is given by
$\mathbf{R}_{x}=\mathbb{E}\big{[}\mathbf{x}[t]\mathbf{x}^{H}[t]\big{]}=\sum_{k\in\mathcal{K}}\mathbf{f}_{k}\mathbf{f}_{k}^{H}+\mathbf{R}_{s}.$ (9)
@@ II-B1 Communication Model
$y_{k}[t]=\underbrace{\mathbf{h}_{k}^{T}\mathbf{f}_{k}c_{k}[t]}_{\text{desired signal}}+\sum_{i\neq k}\mathbf{h}_{k}^{T}\mathbf{f}_{i}c_{i}[t]+\mathbf{h}_{k}^{T}\mathbf{s}[t]+z_{k}[t],$ (10)
@@
$R(\mathbf{f}_{k},\mathbf{R}_{x})=\log_{2}\left(1+\frac{|\mathbf{h}_{k}^{T}\mathbf{f}_{k}|^{2}}{\mathbf{h}_{k}^{T}\mathbf{R}_{x}\mathbf{h}_{k}^{*}-|\mathbf{h}_{k}^{T}\mathbf{f}_{k}|^{2}+\sigma_{k}^{2}}\right).$ (11)
@@ II-B2 Sensing Model
$\mathrm{CRB}(\mathbf{R}_{x},\mathbf{G},\sigma_{s}^{2})=(\mathbf{J}_{11}-\mathbf{J}_{12}\mathbf{J}_{22}^{-1}\mathbf{J}_{12}^{T})^{-1},$ (13)
@@ II-C ISAC Model for Hybrid Digital and Analog Antennas
$\mathbf{x}_{\mathrm{HB}}[t]=\mathbf{F}_{\mathrm{RF}}\big{(}\sum_{k\in\mathcal{K}}\mathbf{f}_{\mathrm{BB},k}c_{k}[t]+\mathbf{s}_{\mathrm{BB}}[t]\big{)},\vspace{-0.2cm}$ (14)
where $\mathbf{F}_{\mathrm{RF}}\in\mathbb{C}^{N\times N_{\mathrm{RF}}}$ denotes the analog beamformer realized by PSs, $\mathbf{f}_{\mathrm{BB},k}\in\mathbb{C}^{N_{\mathrm{RF}}\times 1}$ denotes the digital beamformer for user $k$, and $\mathbf{s}_{\mathrm{BB}}[t]\in\mathbb{C}^{N_{\mathrm{RF}}\times 1}$ denotes the digital dedicated sensing signal, whose covariance is denoted by $\mathbf{R}_{\mathrm{BB},s}$. The covariance matrix of the transmit signal is given by $\mathbf{R}_{\mathrm{HB}}=\mathbf{F}_{\mathrm{RF}}\mathbf{R}_{\mathrm{BB}}\mathbf{F}_{\mathrm{RF}}^{H}$, where $\mathbf{R}_{\mathrm{BB}}=\sum_{k\in\mathcal{K}}\mathbf{f}_{\mathrm{BB},k}\mathbf{f}_{\mathrm{BB},k}^{H}+\mathbf{R}_{\mathrm{BB},s}$. Furthermore, due to the hardware limitation of PSs, the analog beamformer should satisfy a unit-modulus constraint, i.e., $|[\mathbf{F}_{\mathrm{RF}}]_{m,n}|=1,\forall m,n$.
@@ II-C1 Communication Model
The achievable rate achieved by the hybrid digital and analog antenna for user $k$ can be directly obtained according to (11), which is given by $R(\mathbf{f}_{\mathrm{HB},k},\mathbf{R}_{\mathrm{HB}})$, where $\mathbf{f}_{\mathrm{HB},k}=\mathbf{F}_{\mathrm{RF}}\mathbf{f}_{\mathrm{BB},k}$.
@@ III-A Problem Formulation
$\min_{\mathbf{f}_{k},\mathbf{R}_{x}\succeq 0}\quad$ $\mathrm{tr}\big{(}\mathrm{CRB}(\mathbf{R}_{x},\mathbf{G},\sigma_{s}^{2})\big{)},$ (16a)
$\mathrm{s.t.}\quad$ $R(\mathbf{f}_{k},\mathbf{R}_{x})\geq R_{\min,k},\forall k,$ (16b)
@@
$\mathbf{R}_{x}\succeq\sum_{k\in\mathcal{K}}\mathbf{f}_{k}\mathbf{f}_{k}^{H},$ (16d)
where $R_{\min,k}$ denotes the minimum rate requirement of user $k$ and $P_{\max}$ denotes the maximum transmit power. The last constraint (16d) is from the condition that $\mathbf{R}_{s}=\mathbf{R}_{x}-\sum_{k\in\mathcal{K}}\mathbf{f}_{k}\mathbf{f}_{k}^{H}\succeq 0$. The optimization problem for hybrid digital and analog antennas can be formulated similarly.
@@ III-B Proposed Solution for Fully Digital Antennas
We first consider the optimization problem (III-A) for fully digital antennas. One of the main obstacles to solving problem (III-A) is the complex form of the CRB matrix. To solve this issue, we first transform problem (III-A) into the following equivalent but more tractable form [8, Proposition 1]:
$\min_{\mathbf{f}_{k},\mathbf{R}_{x}\succeq 0,\mathbf{U}\succeq 0}\quad$ $\mathrm{tr}(\mathbf{U}^{-1}),$ (17a)
$\mathrm{s.t.}\quad$ $\begin{bmatrix}\mathbf{J}_{11}-\mathbf{U}&\mathbf{J}_{12}\\ \mathbf{J}_{12}^{T}&\mathbf{J}_{22}\end{bmatrix}\succeq 0,$ (17d)
@@
where $\mathbf{U}\in\mathbb{C}^{2\times 2}$ is an auxiliary matrix. In this case, the non-convex objective function of the CRB matrix is transformed into the convex constraint (17d). To address the non-convex constraints (16b) and (16d), the semidefinite relaxation (SDR) is adopted [16]. In particular, define the auxiliary variables $\mathbf{F}_{k}=\mathbf{f}_{k}\mathbf{f}_{k}^{H}$, which satisfies $\mathbf{F}_{k}\succeq 0$ and $\mathrm{rank}(\mathbf{F}_{k})=1$. Then, constraint (16b) can be transformed into the following convex form:
$\gamma_{k}\mathbf{h}_{k}^{T}\mathbf{F}_{k}\mathbf{h}_{k}^{*}\geq\mathbf{h}_{k}^{T}\mathbf{R}_{x}\mathbf{h}_{k}^{*}+\sigma_{k}^{2},$ (18)
@@
$\mathbf{R}_{x}\succeq\sum_{k\in\mathcal{K}}\mathbf{F}_{k}.\vspace{-0.2cm}$ (19)
By omitting the rank-one constraint of $\mathbf{F}_{k}$, the following optimization problem can be obtained:
$\min_{\mathbf{F}_{k}\succeq 0,\mathbf{R}_{x}\succeq 0,\mathbf{U}\succeq 0}\quad$ $\mathrm{tr}(\mathbf{U}^{-1}),$ (20a)
@@
which is convex and can be effectively solved by the standard interior-point algorithm. Although the rank-one constraint is omitted in problem (III-B), given any globally optimal solution $\breve{\mathbf{F}}_{k}$ and $\breve{\mathbf{R}}_{x}$ to problem (III-B), we can always construct the following rank-one solution that achieves the same objective value [6, Theorem 1]:
$\mathbf{f}_{k}^{\star}=(\mathbf{h}_{k}^{T}\breve{\mathbf{F}}_{k}\mathbf{h}_{k}^{*})^{-\frac{1}{2}}\breve{\mathbf{F}}_{k}\mathbf{h}_{k}^{*},\quad\mathbf{R}_{x}^{\star}=\breve{\mathbf{R}}_{x}.$ (21)
Therefore, the above solution must be the globally optimal solution to problem (III-A). With this globally optimal solution, the considered fully digital antenna can provide a theoretical performance upper bound for hybrid digital and analog antennas.
@@ III-C Proposed Solution for Hybrid Digital and Analog Antennas
We continue to study the solution to the optimization problem for hybrid digital and analog antennas. To solve it, we consider a heuristic two-stage optimization framework proposed in [17]. In this framework, the analog beamformer is designed to maximize the array gain at the communication users and the sensing target. Then, the digital beamformers and the digital dedicated sensing signal are optimized with the designed analog beamformer. Let $\mathbf{f}_{\mathrm{RF},\ell}$ denote the $\ell$-th column of $\mathbf{F}_{\mathrm{RF}}$. Then, the analog beamformer can be designed as follows:
$\mathbf{f}_{\mathrm{RF},\ell}=\begin{cases}\mathbf{a}^{*}(r_{\ell},\theta_{\ell}),&1\leq\ell\leq K,\\ \mathbf{a}^{*}(r_{s},\theta_{s}),&K<\ell\leq N_{\mathrm{RF}}.\end{cases}$ (22)
With the above solution of the analog beamformer, the optimization problem is only related to the digital part $\mathbf{f}_{\mathrm{BB},k}$ and $\mathbf{R}_{\mathrm{BB}}$, which has the same form as the problem (III-A). Therefore, it can be optimally solved by the algorithm proposed in the previous section.
@@ IV Numerical Results
In this section, numerical results are provided to verify the effectiveness of the proposed near-field ISAC framework. We assume a BS equipped with a ULA with $N=65$ antennas operating at a frequency of $28$ GHz ($\lambda=1.07$ cm) [10]. The antenna aperture is set to $D=0.5$ m, resulting in a Rayleigh distance of $\frac{2D^{2}}{\lambda}=46.73$ m. There are $K=4$ communication and one sensing target located in the near-field region of the BS. The locations of the users are randomly generated within the near-field region of the BD. The location of the target is set to (20 $(20\text{ m},45^{\circ})$. The maximum transmit power at the BS and the noise power are set to $20$ dBm and $-60$ dBm, respectively. For the hybrid digital and analog antennas, the number of RF chains is set to $N_{\mathrm{RF}}=5$. The minimum rate requirement of each user is assumed to be the same, i.e., $R_{\min,k}=R_{\min},\forall k$. In particular, “FD” represents fully digital antennas, “HB” represents hybrid digital and analog antennas, and “RCRB” represents the root of CRB.

### v5

@@ I Introduction
Furthermore, the near-field effect also provides new possibilities. The spherical wave propagation in the near field introduces a new distance dimension, which has the potential to facilitate joint estimation of distance and angle [9] in sensing and to mitigate interference in communications [10]. To the best of the authors’ knowledge, near-field ISAC systems have not been studied yet, which motivates this work. In this letter, we propose a near-field ISAC framework. Based on the spherical wave propagation, we establish the accurate near-field channel model of the near-field ISAC framework. We jointly optimize the ISAC signal to maximize the near-field sensing performance subject to the minimum near-field communication rate for both fully digital antennas and hybrid digital and analog antennas. Finally, our numerical results verify the effectiveness of the proposed framework.
@@ II System Model
As shown in Fig. 1, we consider a narrowband near-field ISAC system comprising an $N$-antenna dual-functional base station (BS), where $N=2\tilde{N}+1$, $K$ single antenna communication users, and one sensing target. We consider a monostatic sensing setup at the BS. Furthermore, we assume that the BS employs a uniform linear array (ULA) with an antenna spacing of $d$, resulting in an aperture of D=(N1)dD=(N-1)d. Typically, the boundary between near-field and far-field can be determined by the Rayleigh distance $\frac{2D^{2}}{\lambda}$ [3], where $\lambda$ is the signal wavelength. We assume that the communication users and sensing target are located in the near-field region of the BS, which implies that their distance from the BS is less than $\frac{2D^{2}}{\lambda}$. As observed, in order to take advantage of the new features of the near field, ELAAs and high frequencies are required to produce a large near-field region. Therefore, in the following, we first discuss the optimal full digital antennas to provide a performance upper bound of the near-field ISAC system. Then, we extend to the hybrid digital and analog antennas, which are more energy-efficient for ELAAs with high frequencies.
@@ II-A Near-field Channel Model for ISAC
$\mathbf{a}(r,\theta)]_{n}=e^{-j\frac{2\pi}{\lambda}(r_{n}(r,\theta)-r)}.$ (4)
@@
$\mathbf{a}_{\mathrm{far}}(\theta)]_{n}=e^{-j\frac{2\pi}{\lambda}(-nd\cos\theta)}.$ (5)
@@ II-B ISAC Model for Fully Digital Antennas
$\mathbf{x}[t]=\sum_{k\in\mathcal{K}}\mathbf{p}_{k}c_{k}[t]+\mathbf{s}[t],$ (8)
where $\mathbf{p}_{k}\in\mathbb{C}^{N\times 1}$ denotes the fully digital beamformer for conveying the information symbol $c_{k}[t]\in\mathbb{C}$ to the user $k$, $\mathbf{s}[t]$ denotes the dedicated sensing signal for achieving the full sensing degrees of freedom [6], and $\mathcal{K}=\{1,\dots,K\}$. The information symbols are assumed to be independently distributed and have unit power, i.e., $\mathbb{E}\big[c_{k}[t]c_{i}^{*}[t]\big]=1$, if $k=i$; and $\mathbb{E}\big[c_{k}[t]c_{i}^{*}[t]\big]=0$, otherwise. Let $\mathbf{R}_{s}=\mathbb{E}\big[\mathbf{s}[t]\mathbf{s}^{H}[t]\big]$ denote the covariance matrix of the dedicated sensing signal. Then, the covariance matrix of the sensing signal is given by
$\mathbf{R}_{x}=\mathbb{E}\big[\mathbf{x}[t]\mathbf{x}^{H}[t]\big]=\sum_{k\in\mathcal{K}}\mathbf{p}_{k}\mathbf{p}_{k}^{H}+\mathbf{R}_{s}.$ (9)
@@ II-B1 Communication Model
$y_{k}[t]=\underbrace{\mathbf{h}_{k}^{T}\mathbf{p}_{k}c_{k}[t]}_{\text{desired signal}}+\sum_{i\neq k}\mathbf{h}_{k}^{T}\mathbf{p}_{i}c_{i}[t]+\mathbf{h}_{k}^{T}\mathbf{s}[t]+z_{k}[t],$ (10)
@@
$R(\mathbf{p}_{k},\mathbf{R}_{x})=\log_{2}\left(1+\frac{|\mathbf{h}_{k}^{T}\mathbf{p}_{k}|^{2}}{\mathbf{h}_{k}^{T}\mathbf{R}_{x}\mathbf{h}_{k}^{*}-|\mathbf{h}_{k}^{T}\mathbf{p}_{k}|^{2}+\sigma_{k}^{2}}\right).$ (11)
@@ II-B2 Sensing Model
$\mathrm{CRB}(\mathbf{R}_{x},\mathbf{G},\sigma_{s}^{2})=\left(\mathbf{J}_{11}-\mathbf{J}_{12}\mathbf{J}_{22}^{-1}\mathbf{J}_{12}^{T}\right)^{-1},$ (13)
@@ II-C ISAC Model for Hybrid Digital and Analog Antennas
$\mathbf{x}_{\mathrm{HB}}[t]=\mathbf{P}_{\mathrm{RF}}\left(\sum_{k\in\mathcal{K}}\mathbf{p}_{\mathrm{BB},k}c_{k}[t]+\mathbf{s}_{\mathrm{BB}}[t]\right),$ (14)
where $\mathbf{P}_{\mathrm{RF}}\in\mathbb{C}^{N\times N_{\mathrm{RF}}}$ denotes the analog beamformer realized by PSs, $\mathbf{p}_{\mathrm{BB},k}\in\mathbb{C}^{N_{\mathrm{RF}}\times 1}$ denotes the digital beamformer for user $k$, and $\mathbf{s}_{\mathrm{BB}}[t]\in\mathbb{C}^{N_{\mathrm{RF}}\times 1}$ denotes the digital dedicated sensing signal, whose covariance is denoted by $\mathbf{R}_{\mathrm{BB},s}$. The covariance matrix of the transmit signal is given by $\mathbf{R}_{\mathrm{HB}}=\mathbf{P}_{\mathrm{RF}}\mathbf{R}_{\mathrm{BB}}\mathbf{P}_{\mathrm{RF}}^{H}$, where $\mathbf{R}_{\mathrm{BB}}=\sum_{k\in\mathcal{K}}\mathbf{p}_{\mathrm{BB},k}\mathbf{p}_{\mathrm{BB},k}^{H}+\mathbf{R}_{\mathrm{BB},s}$. Furthermore, due to the hardware limitation of PSs, the analog beamformer should satisfy a unit-modulus constraint, i.e., $|[\mathbf{P}_{\mathrm{RF}}]_{m,n}|=1,\forall m,n$.
@@ II-C1 Communication Model
The achievable rate achieved by the hybrid digital and analog antenna for user $k$ can be directly obtained according to (11), which is given by $R(\mathbf{p}_{\mathrm{HB},k},\mathbf{R}_{\mathrm{HB}})$, where $\mathbf{p}_{\mathrm{HB},k}=\mathbf{P}_{\mathrm{RF}}\mathbf{p}_{\mathrm{BB},k}$.
@@ III-A Problem Formulation
$\min_{\mathbf{p}_{k},\mathbf{R}_{x}\succeq 0}\quad$ $\mathrm{tr}\big(\mathrm{CRB}(\mathbf{R}_{x},\mathbf{G},\sigma_{s}^{2})\big),$ (16a)
$\mathrm{s.t.}\quad$ $R(\mathbf{p}_{k},\mathbf{R}_{x})\geq R_{\min,k},\forall k,$ (16b)
@@
$\mathbf{R}_{x}\succeq\sum_{k\in\mathcal{K}}\mathbf{p}_{k}\mathbf{p}_{k}^{H},$ (16d)
where $R_{\min,k}$ denotes the minimum rate requirement of user $k$ and $P_{\max}$ denotes the maximum transmit power. The last constraint (16d) is from the condition that $\mathbf{R}_{s}=\mathbf{R}_{x}-\sum_{k\in\mathcal{K}}\mathbf{p}_{k}\mathbf{p}_{k}^{H}\succeq 0$. The optimization problem for hybrid digital and analog antennas can be formulated similarly.
@@ III-B Proposed Solution for Fully Digital Antennas
We first consider the optimization problem (16) for fully digital antennas. One of the main obstacles to solving problem (16) is the complex form of the CRB matrix. To solve this issue, we first transform problem (16) into the following equivalent but more tractable form [8, Proposition 1]:
$\min_{\mathbf{p}_{k},\mathbf{R}_{x}\succeq 0,\mathbf{U}\succeq 0}\quad$ $\mathrm{tr}(\mathbf{U}^{-1}),$ (17a)
$\mathrm{s.t.}\quad$ $\begin{bmatrix}\mathbf{J}_{11}-\mathbf{U}&\mathbf{J}_{12}\\ \mathbf{J}_{12}^{T}&\mathbf{J}_{22}\end{bmatrix}\succeq 0,$ (17b)
@@
where $\mathbf{U}\in\mathbb{C}^{2\times 2}$ is an auxiliary matrix. In this case, the non-convex objective function of the CRB matrix is transformed into the convex constraint (17b). To address the non-convex constraints (16b) and (16d), the semidefinite relaxation (SDR) is adopted [16]. In particular, define the auxiliary variables $\mathbf{P}_{k}=\mathbf{p}_{k}\mathbf{p}_{k}^{H}$, which satisfies $\mathbf{P}_{k}\succeq 0$ and $\mathrm{rank}(\mathbf{P}_{k})=1$. Then, constraint (16b) can be transformed into the following convex form:
$\gamma_{k}\mathbf{h}_{k}^{T}\mathbf{P}_{k}\mathbf{h}_{k}^{*}\geq\mathbf{h}_{k}^{T}\mathbf{R}_{x}\mathbf{h}_{k}^{*}+\sigma_{k}^{2},$ (18)
@@
$\mathbf{R}_{x}\succeq\sum_{k\in\mathcal{K}}\mathbf{P}_{k}.$ (19)
By omitting the rank-one constraint of $\mathbf{P}_{k}$, the following optimization problem can be obtained:
$\min_{\mathbf{P}_{k}\succeq 0,\mathbf{R}_{x}\succeq 0,\mathbf{U}\succeq 0}\quad$ $\mathrm{tr}(\mathbf{U}^{-1}),$ (20a)
@@
which is convex and can be effectively solved by the standard interior-point algorithm. Although the rank-one constraint is omitted in problem (20), given any globally optimal solution $\widetilde{\mathbf{P}}_{k}$ and $\widetilde{\mathbf{R}}_{x}$ to problem (20), we can always construct the following rank-one solution that achieves the same objective value [6, Theorem 1]:
$\mathbf{p}_{k}^{\star}=(\mathbf{h}_{k}^{T}\widetilde{\mathbf{P}}_{k}\mathbf{h}_{k}^{*})^{-\frac{1}{2}}\widetilde{\mathbf{P}}_{k}\mathbf{h}_{k}^{*},\quad\mathbf{R}_{x}^{\star}=\widetilde{\mathbf{R}}_{x}.$ (21)
Therefore, the above solution must be the globally optimal solution to problem (16). With this globally optimal solution, the considered fully digital antenna can provide a theoretical performance upper bound for hybrid digital and analog antennas.
@@ III-C Proposed Solution for Hybrid Digital and Analog Antennas
We continue to study the solution to the optimization problem for hybrid digital and analog antennas. To solve it, we consider a heuristic two-stage optimization framework proposed in [17]. In this framework, the analog beamformer is designed to maximize the array gain at the communication users and the sensing target. Then, the digital beamformers and the digital dedicated sensing signal are optimized with the designed analog beamformer. Let $\mathbf{p}_{\mathrm{RF},\ell}$ denote the $\ell$-th column of $\mathbf{P}_{\mathrm{RF}}$. Then, the analog beamformer can be designed as follows:
$\mathbf{p}_{\mathrm{RF},\ell}=\begin{cases}\mathbf{a}^{*}(r_{\ell},\theta_{\ell}),&1\leq\ell\leq K,\\ \mathbf{a}^{*}(r_{s},\theta_{s}),&K<\ell\leq N_{\mathrm{RF}}.\end{cases}$ (22)
With the above solution of the analog beamformer, the optimization problem is only related to the digital part $\mathbf{p}_{\mathrm{BB},k}$ and $\mathbf{R}_{\mathrm{BB}}$, which has the same form as the problem (16). Therefore, it can be optimally solved by the algorithm proposed in the previous section.
@@ IV Numerical Results
In this section, numerical results are provided to verify the effectiveness of the proposed near-field ISAC framework. We assume a BS equipped with a ULA with $N=65$ antennas operating at a frequency of $28$ GHz ($\lambda=1.07$ cm) [10]. The antenna aperture is set to $D=0.5$ m, resulting in a Rayleigh distance of $\frac{2D^{2}}{\lambda}=46.73$ m. There are $K=4$ communication and one sensing target located in the near-field region of the BS. The locations of the users are randomly generated within the near-field region of the BD. The location of the target is set to $(20\text{ m},45^{\circ})$. The maximum transmit power at the BS and the noise power are set to $20$ dBm and 60-60 dBm, respectively. For the hybrid digital and analog antennas, the number of RF chains is set to $N_{\mathrm{RF}}=5$. The minimum rate requirement of each user is assumed to be the same, i.e., $R_{\min,k}=R_{\min},\forall k$. In particular, “FD” represents fully digital antennas, “HB” represents hybrid digital and analog antennas, and “RCRB” represents the root of CRB.